# Deblurring 3D Gaussian Splatting

### v1

@@ 3.3 Compensation for Sparse Point Cloud
3D-GS [15] constructs multiple 3D Gaussians from point clouds to model 3D scenes, and its reconstruction quality heavily relies on the initial point clouds. Point clouds are generally obtained from the structure-from-motion (SfM) [36], which extracts features from multi-view images and relates them to several 3D points. However, it can produce only sparse point clouds if the given images are blurry due to the challenge of feature extraction from blurry inputs. Even worse, if the scene has a large depth of field, which is prevalent in defocus blurry scenes, SfM hardly extracts any points that lie on the far end of the scene. To make a dense point cloud, we add additional points after $N_{st}$ iterations. $N_{p}$ points are sampled from an uniform distribution $U(\alpha,\beta)$ where $\alpha$ and $\beta$ are the minimum and maximum value of the position of the points from the existing point cloud, respectively. The color for each new point $p$ is assigned with the interpolated color $p_{c}$ from the nearest neighbors $\mathcal{P}_{\text{knn}}$ among the existing points using K-Nearest-Neigbhor (KNN) [29]. We discard the points whose distance to the nearest neighbor exceeds the distance threshold $t_{d}$ to prevent unnecessary points from being allocated to the empty space. The process of adding extra points to the given point cloud is summarized in algorithm 1. fig. 5 shows that a point cloud with additional points has a dense distribution of points to represent the objects.
Furthermore, 3D-GS [15] effectively manages the number of 3D Gaussians (the number of points) through periodic adaptive density control, densifying and pruning 3D Gaussians. To compensate for the sparsity of 3D Gaussians lying on the far end of the scene, we prune 3D Gaussians depending on their positions. As the benchmark Deblur-NeRF dataset [22] consists of only forward-facing scenes, the z-axis value of each point can be a relative depth from any viewpoints. As shown in fig. 4, we prune out less 3D Gaussians placed on the far edge of the scene to preserve more points located at the far plane, relying on the relative depth. Specifically, the pruning threshold $t_{p}$ is scaled by $w_{p}$ to 1.0 depending on the relative depth, and the lowest threshold is applied to the farthest point. Densifying more for far-plane-3D Gaussians is also an available option, but excessive compensation hinders modeling the scene blurriness, and it can further degrade real-time rendering as extra computational cost is required. We empirically found that flexible pruning was enough to compensate for sparse point cloud, considering the rendering speed. fig. 6 shows that relative depth-based pruning preserves the points lying on the far plane of the scene and leads to successful reconstruction of the objects on the far plane.
Figure 4: Comparison to pruning 3D Gaussians. Left: Given 3D Gaussians. Middle: Applying pruning method proposed by 3D-GS which removes 3D Gaussians with the single threshold ($t_{p}$). Right: Our pruning method which discards unnecessary 3D Gaussians with different thresholds based upon their depth.
@@ 4 Experiments
Table 1: Quantitative results on real defocus dataset. We color each cell as best and second best. *3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
@@
Ours 26.26 0.8240 23.70 0.7268 18.74 0.5539 26.25 0.8197 27.58 0.9037 23.82 0.7430 620
Table 2: Quantitative results on synthetic defocus dataset. We color each cell as best and second best. *3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
@@
Ours 32.03 0.9269 29.89 0.9096 30.50 0.8344 27.56 0.9071 27.18 0.8755 29.43 0.8907 663
We compared our method against the state-of-the-art deblurrig approaches in neural rendering: Deblur-NeRF [22], DP-NeRF [20], PDRF [28] and original 3D Gaussians Splatting (3D-GS) [15]
4.1 Experimental Settings
We utilized PyTorch [27] to create our deblurring pipeline while maintaining the differentiable Gaussian rasterization 3D-GS [15]. We conducted 30k iterations of training. Adam optimizer [16] is used for optimization with the identical setting of 3D-GS. We set the learning rate for MLP at 1e-3, and the rest parameters are identical to 3D-GS. We use MLP with 3 hidden layers, each layer having 64 hidden units and ReLU activation. MLP is initialized with Xavier intialization [11]. For adding extra points to compensate the sparse point cloud, we set the addition start iteration $N_{st}$ to 2,500, the number of supplementing points $N_{p}$ to 100,000, the number of neighbors $K$ to 4, and the minimum distance threshold $t_{d}$ to 10. In terms of depth-based pruning, $t_{p}$ for pruning threshold is set to 5e-3 and 0.3 is used for $w_{p}$, the pruning threshold scaler. All the experiments were conducted on NVIDIA RTX 4090 GPU with 24GB memory.
We evaluated the performance on the benchmark Deblur-NeRF dataset [22] that includes both synthetic and real defocus blurred images. This dataset contains five scenarios in the synthetic defocus category and ten in the real defocus category. Synthetic defocus images are constructed using Blender [5]. By setting the aperture size and randomly selecting a focal plane between the closest and farthest depths, realistic defocus blur effects were produced. COLMAP [36, 37] is used to calculate the camera position of the blur and reference image in the real scene after choosing a big aperture to capture the defocus image.
Figure 5: Comparison on densifying point clouds during training. Left: Example training view. Middle: Point cloud at 5,000 training iterations without adding points. Bottom: Point cloud at 5,000 training iterations with adding extra points at 2,500 iterations.
4.2 Results and Comparisons
First of all, we show the transformation in the values of two Gaussian attributes (scale and rotation) to model the blur. The average values of these two attributes have been mentioned in Table 3. These values have been computed for all the scenes in real defocus dataset. It can be seen that during training, as compared to testing, 3D Gaussians of larger values of scale and rotation have been used to rasterize the scene. This indicates that larger values of scale and rotation are needed to adjust the 3D Gaussians to rasterize the blurred (training) images. While, in contrast, smaller values of these two attributes demonstrate that the smaller sized Gaussians are more suitable to model the fine details that are present in the sharp (testing) images.

### v2

@@ 3.3 Compensation for Sparse Point Cloud
3D-GS [14] constructs multiple 3D Gaussians from point clouds to model 3D scenes, and its reconstruction quality heavily relies on the initial point clouds. Point clouds are generally obtained from the structure-from-motion (SfM) [35], which extracts features from multi-view images and relates them to several 3D points. However, it can produce only sparse point clouds if the given images are blurry. Even worse, if the scene has a large depth of field, which is prevalent in defocus blurry scenes, SfM hardly extracts any points that lie on the far end of the scene. To make a dense point cloud, we add extra points after $N_{st}$ iterations. $N_{p}$ points are sampled from a uniform distribution $U(\alpha,\beta)$ where $\alpha$ and $\beta$ are the minimum and maximum value of the position of the points from the existing point cloud, respectively. The color for each new point $p$ is assigned with the interpolated color $p_{c}$ from the nearest neighbors $\mathcal{P}_{\text{knn}}$ among the existing points using K-Nearest-Neigbhor (KNN) [29]. We discard the points whose distance to the nearest neighbor exceeds the distance threshold $t_{d}$ to prevent unnecessary points from being allocated to the empty space. The process of adding supplementary points to the given point cloud is summarized in Algorithm 1. Fig. 3 shows that a point cloud with additional points has a dense distribution of points to represent the objects.
Figure 3: Comparison on densifying point clouds during training. Left: Example training view. Middle: Point cloud at 5,000 training iterations without adding points. Right: Point cloud at 5,000 training iterations with adding extra points at 2,500 iterations.
Furthermore, 3D-GS [14] effectively manages the number of 3D Gaussians through periodic adaptive density control, densifying and pruning 3D Gaussians. To compensate for the sparsity of 3D Gaussians lying on the far end of the scene, we prune 3D Gaussians depending on their positions. As the benchmark Deblur-NeRF dataset [22] consists of only forward-facing scenes, the z-axis value of each point can be a relative depth from any viewpoint. As shown in Fig. 4, we prune out less 3D Gaussians placed on the far edge of the scene to preserve more points located at the far plane, relying on the relative depth. Specifically, the pruning threshold $t_{p}$ is scaled by $\frac{1}{w_{p}}$ where $w_{p}$ is determined depending on the relative depth, and the lowest threshold is applied to the farthest point.
Figure 4: Comparison to pruning 3D Gaussians. Left: Given 3D Gaussians. Middle: Applying the pruning method proposed by 3D-GS which removes 3D Gaussians with the single threshold ($t_{p}$). Right: Our pruning method that discards unnecessary 3D Gaussians with different thresholds based on their depth.
@@ 4 Experiments
We compared our method against the state-of-the-art deblurring approaches in neural rendering: Deblur-NeRF [22], Sharp-NeRF [18], DP-NeRF [20], PDRF [28] and original 3D Gaussians Splatting (3D-GS) [14] and image-based deblurring which deblurs training images first using Restormer [43] and then trains 3D-GS with them. We evaluated the performance on the benchmark Deblur-NeRF dataset [22] that includes both synthetic and real images captured with either camera motion blur or defocus blur.
4.1 Experimental Settings
We use Adam optimizer [15] and set the learning rate for MLP to $31e-$ 3, that for the position of 3D Gaussians to $31.6e-$ 3. 3D Gaussian pruning threshold ($t_{p}$) and densification threshold are $35e-$ 3 and $42e-$ 4 respectively, for the real defocus blur dataset and $21e-$ 2 and $45e-$ 4 for the real camera motion blur dataset. The rest hyperparameters are identical to those of 3D-GS. We use an MLP with a depth of 4 layers. The first 3 layers are shared for all deltas and features from the shared layers are fed to each of 3 single layer (i.e., 1 layer head for each delta) that produces $\delta x,\delta r$, and $\delta s$ respectively. All layers have 64 hidden units, adopt ReLU activation for non-linearity, and are initialized with Xavier initialization [11]. Both $\lambda_{p}$ and $\lambda_{s}$ are set to $21e-$ 2. For adding extra points to compensate for the sparse point cloud, we set the addition start iteration $N_{st}$ to 2,500, the number of supplementing points $N_{p}$ is proportional to the extent of the point cloud, at most 200,000, further explained in the supplementary material. The number of neighbors $K$ is 4, and the minimum distance threshold $t_{d}$ is 2. In terms of depth-based pruning, the pruning threshold multiplier $w_{p}$ is set to 3. We set $M$ for camera motion deblurring to 5 and the total iteration for training is 20,000. All the experiments were conducted on NVIDIA RTX 4090 GPU.
4.2 Results and Comparisons
As shown in  Tab. 1 and  Fig. 1 our method is on par with the state-of-the-art model in PSNR and achieve state-of-the-art performance evaluated under SSIM on the real defocus blur dataset. Tab. 2 further shows that the proposed method attains state-of-the-art performance on real camera motion blur dataset, under all metrics. At the same time, the proposed method can still enjoy real-time rendering, with a noticeable FPS, while other deblurring models cannot. Fig. 5 shows the qualitative results on real camera motion blur dataset. We can see that ours can produce sharp and fine details, though 3D-GS fails to reconstruct those details. The qualitative and quantitative results on the rest datasets, more experiments and ablation studies are delivered in supplementary materials.
Table 1: Quantitative results on real defocus blur dataset. We color each cell as best and second best.
@@ 4.2 Results and Comparisons
Ours 26.28 0.8235 23.54 0.7310 18.99 0.5705 26.18 0.8166 27.96 0.9098 23.71 0.7471 804
Table 2: Quantitative results on real camera motion blur dataset tested under PSNR, SSIM and FPS. We color each cell as best and second best.
Ball Basket Buick Coffee Decoration
[…]
NeRF [23] 24.08 0.6237 23.72 0.7086 21.59 0.6325 26.47 0.8064 22.39 0.6609
3D-GS [14] 22.99 0.6206 23.11 0.6833 21.22 0.6519 23.53 0.6995 20.45 0.6239
Restormer [43] + 3D-GS 23.85 0.6498 23.75 0.7208 21.42 0.6949 23.94 0.7235 20.98 0.6840
Deblur-NeRF [22] 27.36 0.7656 27.67 0.8449 24.77 0.7700 30.93 0.8981 24.19 0.7707
DP-NeRF [19] 27.20 0.7652 27.74 0.8455 25.70 0.7922 31.19 0.9049 24.31 0.7811
PDRF-10 [28] 27.96 0.7365 28.82 0.8465 25.52 0.7742 31.55 0.8627 23.26 0.7164
Ours 28.27 0.8233 28.42 0.8713 25.95 0.8367 32.84 0.9312 25.87 0.8540
Girl Heron Parterre Puppet Stair Average FPS
[…]
NeRF [23] 20.07 0.7075 20.50 0.5217 23.14 0.6201 22.09 0.6093 22.87 0.4561 22.69 0.6347 $<$ 1
3D-GS [14] 19.72 0.7031 19.26 0.4767 22.22 0.5813 22.18 0.6362 21.88 0.4789 21.66 0.6154 734
Restormer [43] + 3D-GS 19.71 0.7151 19.68 0.5615 22.60 0.6364 22.19 0.6608 22.66 0.5735 22.08 0.6620 708
Deblur-NeRF [22] 22.27 0.7976 22.63 0.6874 25.82 0.7597 25.24 0.7510 25.39 0.6296 25.63 0.7675 $<$ 1
DP-NeRF [19] 23.33 0.8139 22.88 0.6930 25.86 0.7665 25.25 0.7536 25.59 0.6349 25.91 0.7751 $<$ 1
PDRF-10 [28] 23.78 0.8120 22.90 0.6590 25.19 0.7233 25.06 0.7326 25.73 0.5722 25.98 0.7245 $<$ 1
Ours 23.26 0.8390 23.14 0.7438 26.17 0.8144 25.67 0.8051 26.46 0.7050 26.61 0.8224 961
Figure 5: Qualitative results on real camera motion blur dataset.
5 Limitations & Future Works
NeRF-based deblurring methods [22, 20, 28], which are developed under the assumption of volumetric rendering, are not easily applicable to rasterization-based 3D-GS [14]. However, they can be compatible to rasterization by optimizing their MLP to deform kernels in the space of the rasterized image instead of letting MLP deform the rays and kernels in the world space. Although it is an interesting direction, it will incur additional costs for interpolating pixels and just implicitly transform the geometry of 3D Gaussians. Therefore, we believe that it will not be an optimal way to model scene blurriness using 3D-GS [14].
6 Conclusion
We proposed Deblurring 3D-GS, the first deblurring algorithm for 3D-GS. We adopted a small MLP that transforms the 3D Gaussians to model the scene blurriness. We also further facilitated deblurring by complementing more points on sparse point clouds. We validated that our method can deblur the scene while still enjoying the real-time rendering with FPS $>$ 800. This is because we use the MLP only during the training time, and the MLP is not involved in the inference stage, keeping the inference stage identical to the 3D-GS.
Supplementary Materials
Appendix A Additional Method Details
A.0.1 Defocus blur modeling
According to the thin lens law [12], the scene points that lie at the focal distance of the camera make a sharp image at the imaging plane. On the other hand, any scene points that are not at the focal distance will make a blob instead of a sharp point on the imaging plane, and it produces a defocus blurred image. If the separation from the focal distance of a scene point is large, it produces a blob of a large area, which corresponds to severe defocus blur. We assume that 3D Gaussians with greater dispersion (greater scale $s$) can represent the scene points not being located at the focal distance, while those with smaller scale $s$ can model the points placed at the focal distance. Table 3 shows that during training, as compared to testing, 3D Gaussians of larger values of scale have been used to rasterize the scene. This indicates that larger values of scale are needed to adjust the 3D Gaussians to rasterize the blurred (training) images. While, in contrast, smaller values of the attributes demonstrate that the smaller-sized Gaussians are more suitable to model the fine details that are present in the sharp (testing) images.
Table 3: Scale transformation of 3D Gaussians to model the defocus blur measured on real defocus blur dataset.
Train ($\hat{s}_{j}$) Test ($s_{j}$)
Average 0.3232 0.3079
A.0.2 Camera motion blur modeling
Camera motion blur occurs due to the camera shake during the exposure time of a photograph. When a camera moves while capturing an image, the rays from different points within the scene strike different areas of the camera sensor at different moments and are intermixed, so the blurry final image is obtained [7]. Consequently, the blurry final image is acquired as each point in the subject is captured slightly different region on the sensor. Meanwhile, images at all the instants when the rays hit the camera sensor are sharp, without any inherent blur. Therefore, we train the 3D Gaussians to model clean representations at various moments the lights reach the sensor, rasterize $M$ clean images from $M$ different moments, and then average them to simulate the camera movement during training time. Fig. 6 shows the clean images rasterized at different moments and averaged blurry image when $M$ is 6.
Figure 6: Camera motion blur modeling during training. Clean images at different moments of the camera movement during exposure time are rasterized and averaged to a single image to model the camera motion blur.
A.0.3 Selective defocus blurring
The proposed method can handle the training images arbitrarily blurred in various parts of the scene. Since we predict $(\delta s_{j})$ for each Gaussian, we can selectively enlarge the covariances of Gaussians where the parts in the training images are blurred. Such transformed per-Gaussian covariance is projected to 2D space and it can act as pixel-wise blur kernel at the image space. It is noteworthy that applying different shapes of blur kernels to different pixels plays a pivotal role in modeling scene blurriness since blurriness spatially varies. This flexibility enables us to effectively implement deblurring capability in 3D-GS [14]. On the other hand, a naive approach to blurring the rendered image is simply to apply a Gaussian kernel. As shown in Fig. 7, this approach will blur the entire image, not blur pixel-wisely, resulting in blurring parts that should not be blurred for training the model. Even if a learnable Gaussian kernel is applied, optimizing the mean and variance of the Gaussian kernel, a single type of blur kernel is limited in its expressivity to model the complicatedly blurred scene and is optimized to model the average blurriness of the scene from averaging loss function which fails to model blurriness morphing in each pixel. Not surprisingly, the Gaussian blur is a special case of the proposed method. If we predict one $(\delta s_{j})$ for all 3D Gaussians, then it will similarly blur the whole image. Fig. 8 shows that the proposed method successfully deblur the defocus blur, while normal Gaussian blur kernel approaches fail. Moreover, transforming per-Gaussian covariance allows to adjust scene blurriness arbitrarily as shown in Fig. 9.
Figure 7: Comparison to normal Gaussian blur kernel. Top row: It shows the proposed method. $g$ is the Gaussian before the transformation, and $g^{\prime}$ is the Gaussian after the transformation. Since the different transformations can be applied to different Gaussian, ours can selectively blur the images depending on the scene; it can only blur the front parts of the scene. Bottom row: It describes a normal Gaussian blur kernel where $h$ is the Gaussian, and $h^{\prime}$ is the Gaussian after applying a normal Gaussian blur kernel. Simply applying a normal Gaussian blur kernel is not capable of handling different parts of the image differently, thereby uniformly blurring the entire image.
Figure 8: Defocus-deblurred images with different sizes of normal Gaussian blur kernels and the proposed method. (A), (B), (C): 15$\times$15, 9$\times$9, and 5$\times$5 Gaussian blur kernel are in use to deblur, respectively, and the bottom row shows the visualization of each kernel whose values are inverted for better visibility (D): Proposed method which transforms geometry of each 3D Gaussian.
Figure 9: Selective Gaussian blur adjustment. As delineated in  Fig. 7, our methodology adeptly harnesses the $\delta r_{j},\delta s_{j}$ both emanating from compact Multi-Layer Perceptrons (MLP), enabling the inversion of Gaussian blur regions or the comprehensive modulation of overall blurriness and sharpness. With the Transformation of $\delta r_{j},\delta s_{j}$, our framework facilitates the nuanced blurring of proximal regions akin to A, as well as the deft blurring of distant locales akin to B. Furthermore, it offers the capability to manipulate the global blurriness or sharpness, exemplified by adjustments akin to C and D.
A.0.4 Visualization
Fig. 10 visualizes the original and transformed 3D Gaussians for defocus blur. With a given view whose near plane is defocused, the transformed 3D Gaussians show larger scales than those of the original 3D Gaussians to model defocus blur on the near plane (blur-bordered images). Meanwhile, the transformed 3D Gaussians keep very similar shapes to the original ones for sharp objects in the far plane (red-bordered images). Fig. 11 depicts point clouds of the original 3D Gaussians and transformed 3D Gaussians. The point cloud of the transformed 3D Gaussians exhibits the camera movements when it moves left to right.
Figure 10: Gaussians visualization for defocus blur.
Figure 11: Point cloud visualization for camera motion blur.
A.0.5 Adding extra points
The number of supplementing points $N_{p}$ is determined based on the extent of the point cloud as follows:
$N_{p}=\text{min}(\frac{\text{prod}(Q_{max}-Q_{min})}{c^{3}},N_{max}),$ (12)
where $\text{prod}(\cdot)$ returns a product of all values in a vector, $Q_{max}\in\mathbb{R}^{3}$ and $Q_{min}\in\mathbb{R}^{3}$ are the maximum and minimum values of the positions of the points in the point cloud along x-, y-, and z-axis, respectively, and $c$ is a constant for stability which is set to 1.1. $N_{max}$ is 200,000 which restricts the number of points to be added.
Appendix B Additional Experiments
B.0.1 More results
We further evaluate the proposed method under Learned Perceptual Image Patch Similarity (LPIPS) metric as shown in Tab. 4 and Tab. 5 for real defocus blur and real camera motion blur dataset. Our method achieves state-of-the-art performance under LPIPS on both datasets. Also, we conducted experiments on the synthetic defocus blur (Tab. 6, and Tab. 8) and camera motion blur (Tab. 7, and Tab. 8) dataset. For the synthetic defocus blur dataset, we set the learning rate for the position of 3D Gaussians to $44.8e-$ 4. The learning rate for scale and rotation attribute of 3D Gaussians is set to 0.015, and 0.005 respectively, $\lambda_{s}$ to 0.005, and $\lambda_{p}$ to 0.001 for the synthetic camera motion blur dataset. The rest hyperparameters are identical to those for real blur datasets. Qualitative results of each dataset are illustrated in Fig. 14 to Fig. 16.
Table 4: Quantitative results on real defocus blur dataset tested under LPIPS metric. We color each cell as best and second best.
Cake Caps Cisco Coral Cupcake Cups Daisy Sausage Seal Tools Average
NeRF [23] 0.2250 0.2801 0.1256 0.2155 0.2689 0.2315 0.2621 0.2789 0.2680 0.1547 0.2310
3D-GS [14] 0.2082 0.4329 0.1781 0.3111 0.2081 0.3211 0.2629 0.2855 0.3057 0.1953 0.2920
Deblur-NeRF [22] 0.1282 0.1612 0.0868 0.1160 0.1214 0.1271 0.1208 0.1796 0.1048 0.0610 0.1207
DP-NeRF [19] 0.1267 0.1430 0.0840 0.0960 0.1178 0.1035 0.1075 0.1473 0.1026 0.0539 0.1082
PDRF-10 [28] 0.1622 0.2854 0.0943 0.2335 0.1862 0.1370 0.1024 0.2126 0.1927 0.1395 0.1746
Ours 0.1158 0.1491 0.0794 0.1310 0.0985 0.1028 0.0949 0.1405 0.0980 0.0577 0.1068
Table 5: Quantitative results on real camera motion blur dataset tested under LPIPS metric. We color each cell as best and second best.
Ball Basket Buick Coffee Decoration Girl Heron Parterre Puppet Stair Average
NeRF [23] 0.3992 0.3223 0.3502 0.2896 0.3633 0.3196 0.4129 0.4046 0.3389 0.4868 0.3687
3D-GS [14] 0.3505 0.2925 0.2839 0.3113 0.3141 0.2843 0.3328 0.3197 0.2607 0.3888 0.3240
Deblur-NeRF [22] 0.2230 0.1481 0.1752 0.1244 0.1862 0.1687 0.2099 0.2161 0.1577 0.2102 0.1820
DP-NeRF [19] 0.2088 0.1294 0.1405 0.1002 0.1639 0.1498 0.1914 0.1900 0.1505 0.1772 0.1602
PDRF-10 [28] 0.2487 0.1241 0.1751 0.1424 0.2379 0.1828 0.2367 0.2639 0.1569 0.2430 0.2012
Ours 0.1413 0.1155 0.0954 0.0676 0.0933 0.1011 0.1543 0.1206 0.0941 0.1123 0.1096
Table 6: Quantitative results on synthetic defocus blur dataset under PSNR and SSIM metrics. We color each cell as best and second best.
@@ B.0.1 More results
Ours 31.97 0.9275 29.16 0.9089 31.31 0.8580 27.54 0.9083 27.55 0.8858 29.51 0.8977 798
Table 7: Quantitative results on synthetic camera motion blur dataset under PSNR and SSIM metrics. We color each cell as best and second best.
Cozyroom Factory Pool Tanabata Trolley Average FPS
[…]
NeRF [23] 25.66 0.7941 19.32 0.4563 30.45 0.8354 22.22 0.6807 21.25 0.6370 23.78 0.6807 $<$ 1
3D-GS [14] 25.59 0.8076 18.11 0.4179 25.63 0.6326 21.35 0.6686 20.56 0.6257 22.25 0.6305 1019
Deblur-NeRF [22] 32.08 0.9261 25.60 0.7750 31.61 0.8682 27.11 0.8640 27.45 0.8632 28.77 0.8593 $<$ 1
DP-NeRF [19] 32.65 0.9317 25.91 0.7787 31.96 0.8768 27.61 0.8748 28.03 0.8752 29.23 0.8674 $<$ 1
PDRF-10 [28] 31.90 0.9321 26.56 0.8102 31.29 0.8657 28.21 0.8952 28.48 0.8956 29.29 0.8798 $<$ 1
Ours 31.45 0.9222 24.01 0.7333 31.87 0.8829 27.01 0.8807 26.88 0.8710 28.24 0.8580 919
Table 8: Quantiative results on synthetic camera motion blur dataset tested under LPIPS metric. We color each cell as best and second best.
Defocus blur Camera motion blur
Cozy2room Factory Pool Tanabata Wine Average Cozy2room Factory Pool Tanabata Wine Average
NeRF [23] 0.0885 0.2351 0.3340 0.2142 0.2799 0.2303 0.2288 0.5304 0.1932 0.3653 0.3633 0.3362
3D-GS  [14] 0.0692 0.1842 0.5094 0.1710 0.2281 0.2323 0.1645 0.4958 0.2632 0.3235 0.3390 0.3172
Deblur-NeRF [22] 0.0481 0.1127 0.1901 0.0995 0.1436 0.1188 0.0477 0.2687 0.1246 0.1228 0.1363 0.1400
DP-NeRF [19] 0.0386 0.1035 0.1563 0.0779 0.1170 0.0987 0.0355 0.2494 0.0908 0.1033 0.1129 0.1184
PDRF-10 [28] 0.0518 0.1066 0.1893 0.0819 0.1210 0.1101 0.0448 0.1499 0.1345 0.1025 0.0939 0.1051
Ours 0.0424 0.1079 0.1587 0.0821 0.1163 0.1015 0.0367 0.2326 0.0751 0.0785 0.1028 0.1051
B.0.2 Depth-based pruning
We conduct an ablation study on depth-based pruning. To address excessive sparsity of point cloud at the far plane, we prune the points on the far plane less to maintain more numbers of points. Tab. 9 shows our depth-based pruning can preserve more points located at the far plane which leads to better reconstruction quality than naive pruning, which prunes the points with a single threshold regardless of the positions of the points. In addition, Fig. 12 shows a failure to reconstruct objects at the far plane when naive pruning is used, while objects lying on the near-end of the scene are well reconstructed. Meanwhile, ours, with depth-based pruning, can render clean objects on both near and far planes.
Methods PSNR $\uparrow$ SSIM $\uparrow$
Naive pruning 23.52 0.7394
Depth-based pruning 23.73 0.7474